# TFusionOcc: Student’s t-Distribution Based Object-Centric Multi-Sensor Fusion Framework for 3D Occupancy Prediction

### v1

@@ IV-E Model Performance Analysis
To demonstrate the performance of our proposed TFusionOcc, we evaluated it on the Surroundocc-nuScenes and Occ3D-nuScenes benchmarks. For each benchmark, we evaluate our model across three primitive settings, including T-P (general T-primitive), T-SQ (T-superquadric), and T-SQ-IW (deformable T-superquadric or T-superquadric with inverse-warp), with 12800 and 25600 total primitives, respectively. The results of the experiment on the surroundocc-nuScenes and Occ3D-nuScenes benchmarks are shown in Table II and Table III, respectively.
In both benchmarks, TFusionOcc (T-SQ-IW) settings achieve the best and second-best performance, demonstrating the importance of the primitives’ shape variety and robustness. Specifically, our proposed method excels at detecting background objects, including drivable surfaces, other flat areas, sidewalks, terrain, man-made structures, and vegetation. In addition, our model performs significantly well on relatively small foreground objects, such as barriers, bicycles, cars, motorcycles, pedestrians, and traffic cones.
Our method, which only adopts a single 3D supervision signal, achieved par-performance in comparison with algorithms that leverage multi-task learning to receive dual supervision signals, such as DAOcc [33] in the Surroundocc-nuScenes benchmark that incorporates a 3D object detection head and a 3D semantic occupancy prediction head during training, and SDGOcc [4] in the Occ3D-nuScenes benchmark, which incorporates a 2D semantic segmentation head and a 3D semantic occupancy prediction head during training.
@@
Method 2D-Backbone Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [3] R101-DCN C 23.96 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas [20] - C 28.66 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer [13] R101-DCN C 30.50 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer [8] R101-DCN C 30.86 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 C-CONet [29] R101 C 26.10 18.40 18.60 10.00 26.40 27.40 8.60 15.70 13.30 9.70 10.90 20.20 33.00 20.70 21.40 21.80 14.70 21.30 InverseMatrixVT3D [17] R101-DCN C 30.03 18.88 18.39 12.46 26.30 29.11 11.00 15.74 14.78 11.38 13.31 21.61 36.30 19.97 21.26 20.43 11.49 18.47 OccFormer [36] R101-DCN C 31.39 19.03 18.65 10.41 23.92 30.29 10.31 14.19 13.59 10.13 12.49 20.77 38.78 19.79 24.19 22.21 13.48 21.35 FB-Occ [14] R101 C 31.50 19.60 20.60 11.30 26.90 29.80 10.40 13.60 13.70 11.40 11.50 20.60 38.20 21.50 24.60 22.70 14.80 21.60 RenderOcc [22] R101 C 29.20 19.00 19.70 11.20 28.10 28.20 9.80 14.70 11.80 11.90 13.10 20.10 33.20 21.30 22.60 22.30 15.30 20.90 GaussianFormer [9] R101-DCN C 29.83 19.10 19.52 11.26 26.11 29.78 10.47 13.83 12.58 8.67 12.74 21.57 39.63 23.28 24.46 22.99 9.59 19.12 Co-Occ [21] R101 C 30.00 20.30 22.50 11.20 28.60 29.50 9.90 15.80 13.50 8.70 13.60 22.20 34.90 23.10 24.20 24.10 18.00 24.80 GaussianFormer2-256 [7] R101-DCN C 31.74 20.82 21.39 13.44 28.49 30.82 10.92 15.84 13.55 10.53 14.04 22.92 40.61 24.36 26.08 24.27 13.83 21.98 SurroundOcc [30] R101-DCN C 31.49 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 Inverse++ [19] R101-DCN C 31.73 20.91 20.90 13.27 28.40 31.37 11.90 17.76 15.39 13.49 13.32 23.19 39.37 22.85 25.27 23.68 13.43 20.98 LMSCNet [25] - L 36.60 14.90 13.10 4.50 14.70 22.10 12.60 4.20 7.20 7.10 12.20 11.50 26.30 14.30 21.10 15.20 18.50 34.20 L-CONet [29] - L 39.40 17.70 19.20 4.00 15.10 26.90 6.20 3.80 6.80 6.00 14.10 13.10 39.70 19.10 24.00 23.90 25.10 35.70 Co-Occ [21] - L 42.20 22.90 22.00 6.90 25.70 32.40 14.50 13.50 21.00 10.50 18.00 22.50 36.60 21.80 24.60 25.70 31.20 39.90 Co-Occ [21] R101-DCN C+L 41.10 27.10 28.10 16.10 34.00 37.20 17.00 21.60 20.80 15.90 21.90 28.70 42.30 25.40 29.10 28.60 28.20 38.00 GaussianFormer3D [38] R101-DCN C+L 43.30 27.10 26.90 15.80 32.70 36.10 18.60 21.70 24.10 13.00 21.30 29.00 40.60 23.70 27.30 28.20 32.60 42.30 OccFusion [18] R101-DCN C+R 33.97 20.73 20.46 13.98 27.99 31.52 13.68 18.45 15.79 13.05 13.94 23.84 37.85 19.60 22.41 21.20 16.16 21.81 OccFusion [18] R101-DCN C+L 43.53 27.55 25.15 19.87 34.75 36.21 20.03 23.11 25.25 17.50 22.70 30.06 39.47 23.26 25.68 27.57 29.54 40.60 OccFusion [18] R101-DCN C+L+R 43.96 28.27 28.32 20.95 35.06 36.84 20.33 26.22 25.86 19.17 21.27 30.64 40.08 23.75 25.56 27.63 29.82 40.82 GaussianFusionOcc [23] R101-DCN C+L 45.16 30.21 30.22 18.70 35.91 39.57 22.67 27.36 30.10 18.59 24.45 31.25 43.06 25.76 29.12 29.33 34.65 42.70 GaussianFusionOcc [23] R101-DCN C+L+R 45.20 30.37 30.43 18.54 36.23 39.66 22.57 27.35 30.30 19.14 24.56 31.95 42.60 25.82 29.48 29.70 34.78 42.95 OccCylindrical [16] Res-50 C+L 44.94 28.67 26.17 22.12 31.47 36.84 17.95 27.77 29.85 23.90 20.64 28.27 43.00 23.14 27.99 27.81 30.81 40.95 DAOcc [33] Res-50 C+L 45.00 30.50 30.80 19.50 34.00 38.80 25.00 27.70 29.90 22.50 23.20 31.60 41.00 25.90 29.40 29.90 35.20 43.50 TFusionOcc (T-SQ-12800) Res-50 C+L 46.70 30.76 30.67 21.27 33.12 38.96 22.09 27.54 30.47 22.87 21.65 29.86 42.63 27.18 29.80 30.91 36.81 46.26 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 46.92 31.21 31.65 20.13 34.75 39.70 23.08 27.00 30.89 23.33 22.02 30.40 43.11 27.78 30.25 31.47 37.15 46.60 TFusionOcc (T-P-25600) Res-50 C+L 45.49 30.04 30.15 19.58 34.20 38.64 22.10 26.60 29.85 22.24 22.44 29.65 41.08 25.19 28.43 29.71 35.77 45.01 TFusionOcc (T-SQ-25600) Res-50 C+L 46.14 30.88 30.87 21.26 33.68 39.25 22.54 28.34 31.45 23.61 22.80 30.16 41.76 26.12 29.23 30.26 36.77 46.06 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 47.01 31.47 31.96 20.78 34.57 39.73 24.02 27.82 31.19 23.82 23.05 30.67 43.17 27.18 30.13 31.18 37.52 46.71
@@
TABLE II: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation set. Our approach outperforms other existing methods with the same input modality. For readers’ reference, the bottom of the table presents results from three additional methods using different input modalities. All methods are trained with dense occupancy labels from SurroundOcc [30]. Notion of modality: Camera (C), Lidar (L), Radar (R). The best and second-best performance is indicated in bold red and bold blue, respectively.
@@
TABLE III: 3D semantic occupancy prediction results on Occ3D-nuScenes benchmark. All methods are trained with dense occupancy labels from Occ3D [27]. The camera visibility mask is not used during the training phase. Notion of modality: Camera (C), Lidar (L), Radar (R). The best and second-best performance is indicated in bold red and bold blue, respectively.
IV-F Challenging Scenarios Performance Analysis
@@
To further examine the performance and robustness of the proposed algorithm, we follow [18] to evaluate TFusionOcc with different T-primitive representations under challenging rainy and night-time scenarios and the experiment results are shown in Table IV and Table V, respectively.
In the rainy scenario, raindrops on the camera lens cause focus blur and degrade the quality of semantic information provided by the camera branch. Furthermore, laser beam reflections during rainy conditions degraded the quality of the geometry information provided by the lidar branch. As a result, all algorithms in the benchmark exhibit varying degrees of performance degradation. However, our proposed method benefits from a multi-stage feature fusion design and the flexible T-Primitive representation, exhibits the least performance degradation in the benchmark, demonstrating high robustness against challenging rainy conditions.
In the nighttime scenario, due to the nature of the camera sensor, which is sensitive to the ambient lightning conditions, the dim lighting condition during the nighttime leads to semantic information loss in the camera branch. Meanwhile, the depth-estimation accuracy from the camera branch drops. As a result, all algorithms in the benchmark experience severe performance degradation around 1.56% to 11.10% under the IoU metric and 8.89% to 12.50% under the mIoU metric. Among all performance degradation algorithms, our method still achieves SOTA performance, demonstrating superior feature fusion and information filtering capabilities.
Method Backbone Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation InverseMatrixVT3D [17] R101-DCN C 29.72 (-0.31%) 18.99 (+0.11%) 18.55 14.29 22.28 30.02 10.19 15.20 10.03 9.71 13.28 20.98 37.18 23.47 27.74 17.46 10.36 23.13 GaussianFormer [9] R101-DCN C 27.37 (-2.46%) 16.96 (-2.14%) 18.16 9.58 21.09 26.83 8.04 10.13 7.80 5.84 12.66 18.24 35.53 18.51 27.79 19.23 11.04 20.85 Co-Occ [21] R101 C 28.90 (-1.10%) 19.70 (-0.60%) 22.10 17.60 26.30 30.80 10.90 9.90 8.20 9.70 11.40 19.30 39.00 22.20 32.60 23.00 11.50 21.30 GaussianFormer2-256 [7] R101-DCN C 31.14 (-0.60%) 20.36 (-0.46%) 19.84 13.52 26.89 31.65 10.82 15.16 9.04 8.41 13.72 21.84 40.51 24.57 32.21 20.65 12.64 24.33 SurroundOcc [30] R101-DCN C 30.57 (-0.92%) 19.85 (-0.45%) 21.40 12.75 25.49 31.31 11.39 12.65 8.94 9.48 14.51 21.52 35.34 25.32 29.89 18.37 14.44 24.78 Inverse++ [19] R101-DCN C 31.32 (-0.41%) 20.66 (-0.25%) 22.52 13.79 25.49 31.80 11.70 16.72 11.14 10.12 12.29 22.25 38.78 23.93 31.62 21.14 12.65 24.61 Co-Occ [21] R101 C+L 40.30 (-0.80%) 26.60 (-0.50%) 26.60 19.10 37.60 37.20 15.90 20.30 16.30 12.30 23.30 27.00 41.00 22.80 35.20 24.60 27.80 39.30 OccFusion [18] R101-DCN C+L 42.67 (-0.86%) 26.68 (-0.87%) 20.91 18.39 35.26 36.19 17.69 19.05 19.40 17.08 23.86 28.86 38.28 26.37 31.44 21.35 29.48 43.22 OccFusion [18] R101-DCN C+L+R 42.67 (-1.29%) 27.39 (-0.88%) 27.82 21.10 36.00 37.10 17.23 21.67 20.34 17.46 20.93 28.57 38.99 24.72 31.96 21.26 29.64 43.53 GaussianFusionOcc [23] R101-DCN C+L 44.28 (-0.88%) 29.19 (-1.02%) 28.10 19.84 36.28 38.90 18.11 21.13 26.14 17.95 25.79 29.92 41.72 27.35 34.99 22.85 33.84 44.10 GaussianFusionOcc [23] R101-DCN C+L+R 44.36 (-0.84%) 29.86 (-0.51%) 28.40 19.88 38.87 39.33 20.61 26.05 25.66 17.97 26.07 31.02 41.70 24.94 35.27 24.08 33.90 44.00 OccCylindrical [16] Res-50 C+L 44.08 (-0.86%) 28.07 (-0.60%) 24.50 22.66 33.47 36.79 16.96 23.63 23.79 23.49 22.70 28.14 41.67 22.32 33.55 21.44 30.58 43.45 DAOcc [33] Res-50 C+L 44.51 (-0.49%) 29.65 (-0.85%) 28.50 19.63 38.01 38.98 21.42 20.19 24.63 21.52 23.60 31.12 40.56 26.68 35.05 24.94 34.29 45.28 TFusionOcc (T-SQ-12800) Res-50 C+L 45.79 (-0.91%) 30.28 (-0.48%) 29.32 22.06 35.67 38.47 18.43 23.29 26.19 24.65 23.34 28.40 41.79 28.40 34.99 26.16 35.75 47.54 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 45.81 (-1.11%) 30.74 (-0.47%) 30.54 21.85 37.99 39.11 20.43 21.56 26.44 24.82 23.39 28.74 41.99 28.68 35.70 26.86 36.08 47.63 TFusionOcc (T-P-25600) Res-50 C+L 44.77 (-0.72%) 29.22 (-0.82%) 27.86 21.11 36.24 38.33 17.79 18.33 25.25 23.53 25.50 28.59 40.27 24.95 33.54 24.82 34.90 46.57 TFusionOcc (T-SQ-25600) Res-50 C+L 45.40 (-0.74%) 30.38 (-0.50%) 29.18 22.52 36.87 38.61 19.41 22.19 26.98 24.84 24.78 28.70 41.21 27.39 34.74 25.57 35.69 47.44 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 46.25 (-0.76%) 30.82 (-0.65%) 30.38 21.70 38.23 39.28 20.29 19.78 26.75 25.35 24.88 29.05 42.35 28.14 35.91 26.52 36.49 48.01
@@
TABLE IV: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation rainy scenario subset. All methods are trained with dense occupancy labels from SurroundOcc [30]. Notion of modality: Camera (C), Lidar (L), Radar (R). The best and second-best performance is indicated in bold red and bold blue, respectively.
Method Backbone Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation InverseMatrixVT3D [17] R101-DCN C 22.48 (-7.55%) 9.99 (-8.89%) 10.40 12.03 0.00 29.94 0.00 9.92 4.88 0.91 0.00 17.79 29.10 2.37 10.80 9.40 8.68 13.57 GaussianFormer [9] R101-DCN C 20.30 (-9.53%) 9.07 (-10.03%) 6.11 8.70 0.00 25.75 0.00 10.44 2.85 0.55 0.00 17.26 30.65 2.95 12.53 9.94 6.65 10.71 Co-Occ [21] R101 C 18.90 (-11.10%) 9.40 (-10.90%) 4.50 9.30 0.00 29.50 0.00 8.40 3.50 0.00 0.00 15.10 29.40 0.60 12.40 11.50 10.70 15.50 GaussianFormer2-256 [7] R101-DCN C 21.19 (-10.55%) 10.14 (-10.68%) 5.25 9.29 0.00 29.33 0.00 13.65 5.80 0.90 0.00 20.22 31.80 1.94 14.83 10.48 5.96 12.72 SurroundOcc [30] R101-DCN C 24.38 (-7.11%) 10.80 (-9.50%) 10.55 14.60 0.00 31.05 0.00 8.26 5.37 0.58 0.00 18.75 30.72 2.74 12.39 11.53 10.52 15.77 Inverse++ [19] R101-DCN C 23.70 (-8.03%) 10.93 (-9.98%) 8.87 10.19 0.00 32.62 0.00 11.77 7.46 0.72 0.00 22.20 32.95 2.15 13.01 9.79 8.61 14.48 Co-Occ [21] R101 C+L 35.60 (-5.50%) 14.60 (-12.50%) 8.40 16.40 0.00 37.20 0.00 13.80 10.90 0.40 0.00 24.30 36.40 2.20 14.60 17.00 19.90 31.70 OccFusion [18] R101-DCN C+L 40.87 (-2.66%) 15.87 (-11.68%) 13.28 17.53 0.00 36.42 0.00 16.16 11.37 1.42 0.00 25.71 32.64 0.63 16.06 20.87 24.52 37.27 OccFusion [18] R101-DCN C+L+R 41.01 (-2.95%) 16.61 (-11.66%) 15.70 16.26 0.00 38.09 0.00 22.18 13.24 0.08 0.00 25.92 33.15 1.57 16.08 21.09 24.51 37.83 GaussianFusionOcc [23] R101-DCN C+L 42.78 (-2.38%) 18.66 (-11.55%) 16.09 12.27 0.00 39.82 0.00 27.66 13.68 0.07 0.00 38.25 40.10 2.07 19.64 19.82 29.55 39.61 GaussianFusionOcc [23] R101-DCN C+L+R 42.51 (-2.69%) 18.45 (-11.92%) 12.15 11.47 0.00 39.77 0.00 29.58 15.27 0.04 0.00 37.08 37.13 2.58 19.94 20.63 29.42 40.10 OccCylindrical [16] Res-50 C+L 43.38 (-1.56%) 17.79 (-10.88%) 16.19 10.04 0.00 37.84 0.00 25.63 13.28 0.23 0.00 32.93 39.59 3.89 17.89 21.61 26.99 38.47 DAOcc [33] Res-50 C+L 42.85 (-2.15%) 18.53 (-11.97%) 17.18 15.07 0.00 39.05 0.00 25.77 14.70 0.07 0.00 32.51 33.95 4.02 19.78 22.58 30.91 40.87 TFusionOcc (T-SQ-12800) Res-50 C+L 44.74 (-1.96%) 18.81 (-11.95%) 16.01 17.51 0.00 39.84 0.00 23.82 12.76 1.13 0.00 32.79 37.67 3.48 19.70 21.42 31.41 43.40 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 44.82 (-2.10%) 19.07 (-12.14%) 16.81 15.53 0.00 40.36 0.00 21.27 13.33 3.79 0.00 34.89 38.04 3.92 19.86 21.89 31.67 43.72 TFusionOcc (T-P-25600) Res-50 C+L 43.25 (-2.24%) 17.83 (-12.21%) 14.12 10.10 0.00 38.45 0.00 21.99 13.10 1.16 0.00 32.05 34.60 4.61 18.91 23.14 30.76 42.35 TFusionOcc (T-SQ-25600) Res-50 C+L 43.86 (-2.28%) 18.95 (-11.93%) 17.10 16.09 0.00 40.17 0.00 23.21 15.13 2.97 0.00 33.83 36.18 3.75 20.12 20.48 31.30 42.94 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 44.97 (-2.04%) 19.67 (-11.80%) 17.98 19.96 0.00 40.69 0.00 23.30 14.76 3.85 0.00 34.49 38.35 3.43 19.70 22.42 31.95 43.79
TABLE V: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation night scenario subset. All methods are trained with dense occupancy labels from SurroundOcc [30]. Notion of modality: Camera (C), Lidar (L), Radar (R). The best and second-best performance is indicated in bold red and bold blue, respectively.
IV-G Performance Analysis On nuScenes-C Dataset
To further assess the robustness of the proposed algorithm in extreme conditions, such as snow, fog, and sensor malfunctions, we use the nuScenes-C dataset and evaluate TFusionOcc at different levels of corruption and severity. The results of the experiment for several camera corruption settings are shown in Tables VI, VII, and X, and the results for several lidar corruption settings are shown in Tables VIII, IX, and XI. Note that since the nuScenes-C dataset only provides single-lidar-sweep corruption data, and our model leverages 10 lidar sweeps. Hence, when evaluating our model across different lidar corruption settings, we use only a single corrupted lidar sweep from the nuScenes-C dataset, and the model performance under the clean setting differs from that in Table II due to a much sparser lidar point cloud. In the camera corruption scenario, our model shows strong robustness against camera crashes, frame loss, brightness, and fog, as measured by the mIoU metric. Nevertheless, the motion blur and snow scenarios cause significant performance degradation in our model. With around 5.04% to 8.43% and 4.91% to 6.64% degradation under the IoU metric, 10.00% to 10.63% and 14.02% to 14.55% degradation under the mIoU metric for the motion blur and snow corruption case, respectively. In the lidar corruption scenario, our model is sensitive to lidar point cloud density and quality degradation, demonstrating its firm reliance on lidar point cloud quality.

### v2

@@ IV-D Model Performance Analysis
To demonstrate the performance of our proposed TFusionOcc, we evaluated it on the SurroundOcc-nuScenes and Occ3D-nuScenes benchmarks. For each benchmark, we evaluate our model across three primitive settings, including T-P (plain T-primitive), T-SQ (T-superquadric), and T-SQ-IW (deformable T-superquadric or T-superquadric with inverse-warp), with 12800 and 25600 total primitives, respectively. The results of the experiment on the SurroundOcc-nuScenes and Occ3D-nuScenes benchmarks are shown in Table II and Table III, respectively.
@@
Method 2D-Backbone Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation MonoScene [3] R101-DCN C 23.96 7.31 4.03 0.35 8.00 8.04 2.90 0.28 1.16 0.67 4.01 4.35 27.72 5.20 15.13 11.29 9.03 14.86 Atlas [20] - C 28.66 15.00 10.64 5.68 19.66 24.94 8.90 8.84 6.47 3.28 10.42 16.21 34.86 15.46 21.89 20.95 11.21 20.54 BEVFormer [13] R101-DCN C 30.50 16.75 14.22 6.58 23.46 28.28 8.66 10.77 6.64 4.05 11.20 17.78 37.28 18.00 22.88 22.17 13.80 22.21 TPVFormer [8] R101-DCN C 30.86 17.10 15.96 5.31 23.86 27.32 9.79 8.74 7.09 5.20 10.97 19.22 38.87 21.25 24.26 23.15 11.73 20.81 C-CONet [30] R101 C 26.10 18.40 18.60 10.00 26.40 27.40 8.60 15.70 13.30 9.70 10.90 20.20 33.00 20.70 21.40 21.80 14.70 21.30 InverseMatrixVT3D [17] R101-DCN C 30.03 18.88 18.39 12.46 26.30 29.11 11.00 15.74 14.78 11.38 13.31 21.61 36.30 19.97 21.26 20.43 11.49 18.47 OccFormer [38] R101-DCN C 31.39 19.03 18.65 10.41 23.92 30.29 10.31 14.19 13.59 10.13 12.49 20.77 38.78 19.79 24.19 22.21 13.48 21.35 FB-Occ [14] R101 C 31.50 19.60 20.60 11.30 26.90 29.80 10.40 13.60 13.70 11.40 11.50 20.60 38.20 21.50 24.60 22.70 14.80 21.60 RenderOcc [22] R101 C 29.20 19.00 19.70 11.20 28.10 28.20 9.80 14.70 11.80 11.90 13.10 20.10 33.20 21.30 22.60 22.30 15.30 20.90 GaussianFormer [9] R101-DCN C 29.83 19.10 19.52 11.26 26.11 29.78 10.47 13.83 12.58 8.67 12.74 21.57 39.63 23.28 24.46 22.99 9.59 19.12 Co-Occ [21] R101 C 30.00 20.30 22.50 11.20 28.60 29.50 9.90 15.80 13.50 8.70 13.60 22.20 34.90 23.10 24.20 24.10 18.00 24.80 GaussianFormer2-256 [7] R101-DCN C 31.74 20.82 21.39 13.44 28.49 30.82 10.92 15.84 13.55 10.53 14.04 22.92 40.61 24.36 26.08 24.27 13.83 21.98 QuadricFormer [42] R101-DCN C 31.22 20.12 19.58 13.11 27.27 29.64 11.25 16.26 12.65 9.15 12.51 21.24 40.20 24.34 25.69 24.24 12.95 21.86 SurroundOcc [31] R101-DCN C 31.49 20.30 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86 Inverse++ [19] R101-DCN C 31.73 20.91 20.90 13.27 28.40 31.37 11.90 17.76 15.39 13.49 13.32 23.19 39.37 22.85 25.27 23.68 13.43 20.98 LMSCNet [25] - L 36.60 14.90 13.10 4.50 14.70 22.10 12.60 4.20 7.20 7.10 12.20 11.50 26.30 14.30 21.10 15.20 18.50 34.20 L-CONet [30] - L 39.40 17.70 19.20 4.00 15.10 26.90 6.20 3.80 6.80 6.00 14.10 13.10 39.70 19.10 24.00 23.90 25.10 35.70 Co-Occ [21] - L 42.20 22.90 22.00 6.90 25.70 32.40 14.50 13.50 21.00 10.50 18.00 22.50 36.60 21.80 24.60 25.70 31.20 39.90 Co-Occ [21] R101-DCN C+L 41.10 27.10 28.10 16.10 34.00 37.20 17.00 21.60 20.80 15.90 21.90 28.70 42.30 25.40 29.10 28.60 28.20 38.00 GaussianFormer3D [39] R101-DCN C+L 43.30 27.10 26.90 15.80 32.70 36.10 18.60 21.70 24.10 13.00 21.30 29.00 40.60 23.70 27.30 28.20 32.60 42.30 OccFusion [18] R101-DCN C+R 33.97 20.73 20.46 13.98 27.99 31.52 13.68 18.45 15.79 13.05 13.94 23.84 37.85 19.60 22.41 21.20 16.16 21.81 OccFusion [18] R101-DCN C+L 43.53 27.55 25.15 19.87 34.75 36.21 20.03 23.11 25.25 17.50 22.70 30.06 39.47 23.26 25.68 27.57 29.54 40.60 OccFusion [18] R101-DCN C+L+R 43.96 28.27 28.32 20.95 35.06 36.84 20.33 26.22 25.86 19.17 21.27 30.64 40.08 23.75 25.56 27.63 29.82 40.82 GaussianFusionOcc [23] R101-DCN C+L 45.16 30.21 30.22 18.70 35.91 39.57 22.67 27.36 30.10 18.59 24.45 31.25 43.06 25.76 29.12 29.33 34.65 42.70 GaussianFusionOcc [23] R101-DCN C+L+R 45.20 30.37 30.43 18.54 36.23 39.66 22.57 27.35 30.30 19.14 24.56 31.95 42.60 25.82 29.48 29.70 34.78 42.95 OccCylindrical [16] Res-50 C+L 44.94 28.67 26.17 22.12 31.47 36.84 17.95 27.77 29.85 23.90 20.64 28.27 43.00 23.14 27.99 27.81 30.81 40.95 DAOcc [35] Res-50 C+L 45.00 30.50 30.80 19.50 34.00 38.80 25.00 27.70 29.90 22.50 23.20 31.60 41.00 25.90 29.40 29.90 35.20 43.50 TFusionOcc (T-SQ-12800) Res-50 C+L 46.70 30.76 30.67 21.27 33.12 38.96 22.09 27.54 30.47 22.87 21.65 29.86 42.63 27.18 29.80 30.91 36.81 46.26 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 46.92 31.21 31.65 20.13 34.75 39.70 23.08 27.00 30.89 23.33 22.02 30.40 43.11 27.78 30.25 31.47 37.15 46.60 TFusionOcc (T-P-25600) Res-50 C+L 45.49 30.04 30.15 19.58 34.20 38.64 22.10 26.60 29.85 22.24 22.44 29.65 41.08 25.19 28.43 29.71 35.77 45.01 TFusionOcc (T-SQ-25600) Res-50 C+L 46.14 30.88 30.87 21.26 33.68 39.25 22.54 28.34 31.45 23.61 22.80 30.16 41.76 26.12 29.23 30.26 36.77 46.06 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 47.01 31.47 31.96 20.78 34.57 39.73 24.02 27.82 31.19 23.82 23.05 30.67 43.17 27.18 30.13 31.18 37.52 46.71
@@
TABLE II: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation set. Notion of modality: Camera (C), LiDAR (L), Radar (R). The best and second-best performance is indicated in bold red and bold blue, respectively.
@@
TABLE III: 3D semantic occupancy prediction results on Occ3D-nuScenes benchmark. Notion of modality: Camera (C), LiDAR (L), Radar (R). The best and second-best performance is indicated in bold red and bold blue, respectively.
In both benchmarks, TFusionOcc (T-SQ-IW) settings achieve the best and second-best performance, demonstrating the importance of the primitives’ shape variety and robustness. Specifically, our proposed method excels at detecting background objects, including drivable surfaces, other flat areas, sidewalks, terrain, man-made structures, and vegetation. In addition, our model performs significantly well on relatively small foreground objects, such as barriers, bicycles, cars, motorcycles, pedestrians, and traffic cones. Notably, our method, which only adopts a single 3D supervision signal, achieved par-performance in comparison with algorithms that leverage multi-task learning to receive dual supervision signals, such as DAOcc [35] in the SurroundOcc-nuScenes benchmark that incorporates an auxiliary 3D object detection head to leverage an extra 3D supervision signal during training, and SDGOcc [4] in the Occ3D-nuScenes benchmark, which incorporates an auxiliary 2D semantic segmentation head to utilize an extra 2D supervision signal during training. It is also worth noting that, as highlighted in the green and red rows in both benchmarks, for the same total number of primitive settings, the primitive with greater shape variety achieves better performance. In addition, increasing the number of primitives from 12,800 to 25,600 usually yields further gains, though the improvement is moderate compared with that from stronger primitive geometry. This suggests that primitive shape flexibility is a more decisive factor than simply increasing primitive count.
@@
To further examine the performance and robustness of the proposed algorithm, we follow [18] to evaluate TFusionOcc in settings under challenging rainy and night-time scenarios, and the experiment results are shown in Table IV.
Method Backbone Modality Rainy IoU / mIoU Night IoU / mIoU Co-Occ [21] R101 C+L 40.30 / 26.60 35.60 / 14.60 OccFusion [18] R101-DCN C+L+R 42.67 / 27.39 41.01 / 16.61 GaussianFusionOcc [23] R101-DCN C+L 44.28 / 29.19 42.78 / 18.66 GaussianFusionOcc [23] R101-DCN C+L+R 44.36 / 29.86 42.51 / 18.45 OccCylindrical [16] Res-50 C+L 44.08 / 28.07 43.38 / 17.79 DAOcc [35] Res-50 C+L 44.51 / 29.65 42.85 / 18.53 TFusionOcc (T-SQ-12800) Res-50 C+L 45.79 / 30.28 44.74 / 18.81 TFusionOcc (T-SQ-IW-12800) Res-50 C+L 45.81 / 30.74 44.82 / 19.07 TFusionOcc (T-P-25600) Res-50 C+L 44.77 / 29.22 43.25 / 17.83 TFusionOcc (T-SQ-25600) Res-50 C+L 45.40 / 30.38 43.86 / 18.95 TFusionOcc (T-SQ-IW-25600) Res-50 C+L 46.25 / 30.82 44.97 / 19.67
@@
TABLE IV: 3D semantic occupancy prediction results on SurroundOcc-nuScenes validation rainy and nighttime scenario subset. The best and second-best performance is indicated in bold red and bold blue, respectively.
In both challenging scenarios, our proposed method under the T-SQ-IW setting achieved the best and second-best performance, demonstrating its robustness and the importance of the geometric shape representation capability of primitives.
IV-E Performance Analysis on nuScenes-C Dataset
To thoroughly assess the robustness of the proposed algorithm under extreme conditions, such as snow, fog, and sensor malfunctions, we use the nuScenes-C dataset and evaluate TFusionOcc across varying levels of corruption and severity. The results of the experiment for several camera corruption settings are shown in Table V.